OpenStudy (anonymous):

find f(g(x)) and g(f(x)) when f(x)=6x-7 and g(x)=6x+7

OpenStudy (anonymous):

f(g(x)) = 6(6x+7) -7 = 36x- 35

myininaya (myininaya):

f(g(x))=f(6x+7)=6(6x+7)-7 g(f(x))=g(6x-7)=6(6x-7)+7

OpenStudy (anonymous):

faultless algebra myin!

myininaya (myininaya):

no u lol